OpenStudy (anonymous):

A rectangular storage container with an open top is to have a volume of 26 cubic meters. The length of its base is twice the width. Material for the base costs 12 dollars per square meter. Material for the sides costs 9 dollars per square meter. Find the cost of materials for the cheapest such container

OpenStudy (surry99):

Can you write the equation for the cost function?

OpenStudy (anonymous):

l = length h = height w = width v = volume Given: v = 40 l = 2w Substituting v=40 and l=2w into volume formula: V = l * w * h 40 = 2w * w * h 40 = 2w^2 * h Solving for h: 40/2w^2 = h 20/w^2 = h Determine the cost function: Cost = 2*(end area)*$6 + 2*(side area)*$6 + (bottom area)*$40 Cost = 12*(h*w) + 12*(h*l) + 40*(w*l) Substitute l=2w: Cost = 12hw + 12*(h*2w) + 40*(w*2w) Cost = 12hw + 24hw + 80w^2 Cost = 36hw + 80w^2 Substitute h=20/w^2: Cost = 36(20/w^2)w + 80w^2 Cost = 36(20/w) + 80w^2 Cost = 720/w + 80w^2 To find the min/max, take the derivative: Cost' = -720/w^2 + 160w and find the zeros: 0 = -720/w^2 + 160w 720/w^2 = 160w 720/160 = w^3 4.5 = w^3 1.65096362 = w Using l=2w: l = 2* 1.65096362 l = 3.301927249 Using h=20/w^2: h = 20/(1.65096362)^2 h = 20/(1.65096362)^2 h = 20/2.7256808745635044 h = 7.337616148 If we take the second derivative: Cost' = -720w^(-2) + 160w Cost'' = 1440w^(-3) + 160 At the point w=1.65096362, the function is concave up, thus this is a local minimum. To determine the cost at this point: Cost = 720/w + 80w^2 Cost = 720/(1.65096362) + 80*(1.65096362)^2 Cost = 654.16 Therefore, the solution is: Width = 1.65096362 Length = 3.301927249 Height = 7.337616148 Cost = 654.16

OpenStudy (anonymous):

holy cow... Thank you! This problem has been giving me trouble for almost 45 minutes now

OpenStudy (surry99):

unfortunately , the answer they provided is wrong...compare the variables in your problems to theirs...For example your volume = 26 m^3....they used 40

OpenStudy (anonymous):

O I see, ya they used different variables... At least I know how to solve it now haha

OpenStudy (anonymous):

medal guys

OpenStudy (anonymous):

be nice